OpenStudy (anonymous):

The Richter scale converts seismographic readings into numbers for measuring the magnitude of an earthquake according to this function M(x) = log , where x0 = 10-3. Two earthquakes differ by 0.1 when measured on the Richter scale. How would the seismographic readings differ at a distance of 100 kilometers from the epicenter?

hartnn (hartnn):

please give the function M(x) properly.

OpenStudy (anonymous):

m(x)=log(x/x0)

OpenStudy (anonymous):

x0=10^-3

hartnn (hartnn):

ok,so let the 2 readings be x1 and x2 So M(x1)-M(x2)=0.1 so log(x1/x0)-log(x2/x0)=0.1 ok till here ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so using log properties: log A-log B =log (A/B) we get log(x1/x2)= 0.1 (x0 gets cancelled) ok? so x1/x2 = 10^(0.1)

hartnn (hartnn):

so using log properties: log A-log B =log (A/B) we get log(x1/x2)= 0.1 (x0 gets cancelled) ok? so x1/x2 = 10^(0.1)

OpenStudy (anonymous):

yessss i get it :P

hartnn (hartnn):

so 10^0.1=1.259 so they will differ by 1.259.